If Steven can mix 20 drinks in 5 minutes, Sue can mix 20 drinks in 10 minutes, and Jack can mix 20 drinks in 15 minutes, how much time will it take all 3 of them working together to mix the 20 drinks?
Answer by ReadingBoosters(3246)

Rate(Steven): 20 / 5 min
Rate(Sue): 20 / 10 min
Rate(Jack): 20 / 15 min
Combined rate:
20/5 + 20/10 + 20/15 = 20/x where x is their combined rate
Divide both sides by 20
1/5 + 1/10 + 1/15 = 1/x
Create common denominators
6/30 + 3/30 + 2/30 = 1/x
11/30 = 1/x
Knowing 1/x is the inverse of x, invert the other side as well
x = 30/11 = 2 8 /11 = 2.72 minutes
.........................
Delighted to help.
-Reading Boosters
Wanting for others what we want for ourselves.
www.MyHomeworkAnswers.com